undeadknight26 (undeadknight26):

Which polynomial identity will prove that 16 = 25 - 9? Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial

undeadknight26 (undeadknight26):

Pretty sure its B...just making sure...

OpenStudy (anonymous):

16 = 25 - 9 (4)^2 = (5)^2 - (3)^2 so its A